The Two Dollar Store has a cost of equity of 10.8 percent, the YTM on the company's bonds is 6.1 percent, and the tax rate is 35 percent. If the company's debt–equity ratio is .55, what is the weighted average cost of capital?
Answer #1

Weight of equity = 1 / (1 + 0.55) = 0.6452

Weight of debt = 1 - 0.6452 = 0.3548

WACC = Weight of debt*after tax cost of debt + weight of equity*cost of equity

WACC = 0.3548*0.061*(1 - 0.35) + 0.6452*0.108

WACC = 0.01407 + 0.06968

WACC = 0.0838 or 8.38%